# Fluctuations of blowup time in a simple model of a super-Malthusian catastrophe

### v1

@@ I Introduction
$\dot{n}(t)=\beta n^{2}(t)\,,\quad\beta>0\,,$ (1)
@@
$n(t)=\frac{n_{0}}{1-\beta n_{0}t},$ (2)
where $n_{0}\equiv n(t=0)>0$ is the initial condition, blows up in a finite time $T=(\beta n_{0})^{-1}$.
@@
$\dot{x}(t)=\beta x^{2}(t)+\eta(t)\,,$ (3)
where $\eta(t)$ is white Gaussian noise, and we have replaced $n$ by $x$. The finite-time blowup statistics for this model and its extensions has been recently studied in Ref. Krapivsky and Meerson (2025), see also Refs. Ornigotti et al. (2018); Šiler et al. (2018); Ryabov et al. (2019).
@@
where $P_{n}(t)$ is the probability of observing $n$ particles at time $t$. We suppose that there are $m\geq 2$ particles at $t=0$, so the initial condition is
@@ II.1 Average blowup time
In particular, for $m=2$ (there are exactly two particles at $t=0$) we obtain $\Theta(2)\equiv\langle T\rangle=2$. It is twice as large as the mean-field result $t_{0}=1$ for the same initial number of particles. Notice that the average blowup time (8) falls off quite slowly as a function of $m$.
@@ II.3 Short-time tail and optimal trajectory
This result is asymptotically exact at $T\to 0$, when the action $S_{0}$ tends to infinity. As we can see, Eq. (41) agrees in the leading order with the short-time tail (28) that we obtained from the exact solution. It misses, however, the important pre-exponential factor $\sim T^{-7/2}\gg 1$ which appears in Eq. (28). The calculation of this factor with the OFM formalism would demand going to the subleading order, which we do not pursue here.

### v3

@@ I Introduction
$\dot{n}(t)=\alpha n^{2}(t)\,,\quad\alpha>0\,,$ (1)
@@
$n(t)=\frac{n_{0}}{1-\alpha n_{0}t},$ (2)
where $n_{0}\equiv n(t=0)>0$ is the initial condition, blows up in a finite time $T=(\alpha n_{0})^{-1}$.
@@
$\dot{x}(t)=\alpha x^{2}(t)+\eta(t)\,,$ (3)
where $\eta(t)$ is white Gaussian noise, and we have replaced $n$ by $x$. The finite-time blowup statistics for this model and its extensions has been recently studied in Ref. Krapivsky and Meerson (2025), see also Refs. Ryabov (2016); Ornigotti et al. (2018); Šiler et al. (2018); Ryabov et al. (2019).
@@
where $P_{n}(t)$ is the probability of observing $n$ particles at time $t$, and $\beta=2\alpha$. We suppose that there are $m\geq 2$ particles at $t=0$, so the initial condition is
@@ II.1 Average blowup time
In particular, for $m=2$ (there are exactly two particles at $t=0$) we obtain $\Theta(2)\equiv\langle T\rangle=2$. It is twice as large as the mean-field prediction $t_{0}=1$ for the same initial number of particles. For $m\gg 1$ the mean-field prediction agrees with the exact result as to be expected. Notice that the average blowup time (8) falls off quite slowly as a function of $m$.
@@ II.3 Short-time tail and optimal trajectory
This result is asymptotically exact at $T\to 0$, when the action $S_{0}$ tends to infinity. As we can see, Eq. (41) agrees in the leading order with the short-time tail (28) that we obtained from the exact solution. It misses, however, the important pre-exponential factor $\sim T^{-7/2}\ll 1$ which appears in Eq. (28). The calculation of this factor in the OFM formalism would demand going to the subleading order, which we do not pursue here.